OpenStudy (anonymous):

Jeremy wants to buy a new computer. The sales woman says that he can make a down payment and then pay for the ocmputer in stallments. Here's a formula that describes this scenario: t=x+yz where t=total price x=amount down y=money each month z=number of months. If Jeremy puts $200 down and pays $50.00 per month for 24 months, how much will the computer cost ?

OpenStudy (anonymous):

@31356

OpenStudy (31356):

X=t-yz 1.Add yz both sides. X+yz=t-yz+yz X+yz=t 2.Subtract X both sides. X+yz-X=t-X yz=t-X

OpenStudy (31356):

y=t-x/z is the answer

OpenStudy (31356):

Medal?

OpenStudy (31356):

@JC9977002

OpenStudy (anonymous):

Of course (: thank you

OpenStudy (31356):

Your welcome

OpenStudy (anonymous):

@31356 it has to be a number lol

OpenStudy (the_fizicx99):

No its not. t=x+yz where t=total price x=amount down y=money each month z=number of months. If Jeremy puts $200 down and pays $50.00 per month for 24 months, how much will the computer cost ? Basically plug in, your solving for t. t = 200 + 50(24) t = 200 + 1200 t = 1400 The total amount of the computer is $1,400

OpenStudy (31356):

Oh yeah forgot to do the other part

OpenStudy (the_fizicx99):

._. ^

OpenStudy (anonymous):

thank you (:

OpenStudy (31356):

ust direct substitution. t = 200 + 50(24) t = $1,400.00

OpenStudy (the_fizicx99):

I don't know whatever is you did @31356 but that's wrong.

OpenStudy (31356):

Actually use direct subsitiution

OpenStudy (anonymous):

posting another one in a minute

OpenStudy (31356):

Sorry yeah that's wrong I miscalcualted

OpenStudy (31356):

To put your formula in detail, (total price) = (downpayment) + (monthly payment)(number of months) where downpayment = $200 monthly payment = $50 number of months = 24 months The variable to be solved is the total price. Using the detailed formula above, total price = (200) + (50)(24) total price = 200 + 1200 total price = 1400 Therefore, to answer your question, Jeremy will pay a total of $1400 after 24 months.

OpenStudy (31356):

Now this is right, sorry

OpenStudy (31356):

@JC9977002 , medal?

OpenStudy (the_fizicx99):

Did you seriously just copy and paste that from yahoo?

OpenStudy (anonymous):

who gets the medal ._.

OpenStudy (the_fizicx99):

I don't really care. 31356, you're violation the CoC and ToA of this website by copying and pasting and taking credit for it. Pretty much plagiarism. @e.mccormick

OpenStudy (31356):

No, there is one in Yahoo I didn't know that

OpenStudy (31356):

I didn't do that, I just did it from scratch.

OpenStudy (31356):

Why would I do that I mean, seriously?